Solve by HAND.

Initial Cost : $4,200

Maintenance every 4 years: $250

Annual Cost: $0

Salvage Value: $40

i = 10%

n = 20 years

Calculate PW,AW,FW
Answer #1

Here annual cost is given 0

PW = -4200 - 250*[(P/F,10%,4) + (P/F,10%,8) + (P/F,10%,12) + (P/F,10%,16)] + 40*(P/F,10%,20)

= -4200 - 250*[0.683013 + 0.466507 + 0.318631 + 0.217629] + 40*0.148644

= -4612.50

FW = -4612.50 * (F/P,10%,20) = -4612.50 * 6.7275 = -31050.77

AW = -4612.50 * (A/P,10%,20) = -4612.50 * 0.117460 = -542.13